OpenStudy (anonymous):

Translate this problem to an equation and solve.If a number is subtracted from its square, the result is 2. Find the number. @Noura11

OpenStudy (anonymous):

Let be x that number ! Then we can write : \[x^2-x=2\] then : \[x^2-x-2=0\] So : \[\Delta=(-1)^2-4\times(-2)=9\] So : \[x_1=\frac{1+\sqrt9}{2}=4\\ x_2=\frac{1-\sqrt9}{2}=-1 \] So there are two numbers verifies the problem !